# An Information-Theoretic Perspective on Multi-LLM Uncertainty Estimation

### v1

@@ Abstract
Large language models (LLMs) often behave inconsistently across inputs, indicating uncertainty and motivating the need for its quantification in high-stakes settings. Prior work on calibration and uncertainty quantification often focuses on individual models, overlooking the potential of model diversity. We hypothesize that LLMs make complementary predictions due to differences in training and the Zipfian nature of language, and that aggregating their outputs leads to more reliable uncertainty estimates. To leverage this, we propose MUSE (Multi-LLM Uncertainty via Subset Ensembles), a simple information-theoretic method that uses Jensen-Shannon Divergence to identify and aggregate well-calibrated subsets of LLMs. Experiments on binary prediction tasks demonstrate improved calibration and predictive performance compared to single-model and naïve ensemble baselines.
An Information-Theoretic Perspective on Multi-LLM Uncertainty Estimation
Maya Kruse1, Majid Afshar3, Saksham Khatwani1,2, Anoop Mayampurath3, Guanhua Chen3 Yanjun Gao1 1University of Colorado Anschutz Medical Campus 2University of Colorado Boulder 3University of Wisconsin Madison Correspondence: yanjun.gao@cuanschutz.edu
@@ 3 LLM Uncertainty Quantification
To estimate uncertainty, we apply a bootstrapping procedure: we resample $90\%$ of the outputs with replacement and recompute $\hat{p}^{\texttt{yes}}$ over $B=100$ trials (denoting as GENBS). From the resulting $\hat{p}^{\texttt{yes}}$ distribution, we compute variance, entropy, and JSD for our proposed algorithms (§ 4).
(2) Sequence Likelihood Scoring. We adopt the sequence likelihood (SLL) approach used in prior LLM calibration work. For each input $x$, we compute the total log-likelihood of two candidate completions: "Answer is Yes" and "Answer is No", denoted $\text{LL}_{\texttt{yes}}$ and $\text{LL}_{\texttt{no}}$. These are computed using left-to-right autoregressive decoding: $\text{LL}_{\texttt{label}}=\sum_{t=1}^{T}\log P(y_{t}^{\texttt{label}}\mid x,y_{<t}^{\texttt{label}})$. The final prediction probability is obtained via softmax normalization. We use this predicted distribution to compute metrics such as AUROC, ECE and Brier Scores Guo et al. (2017). Although not robust to LLM output variability, SLL provides a deterministic scoring baseline for comparison.
4 Multi-LLM selective algorithm
When multiple LLMs generate predictive distributions over binary outcomes, we interpret the disagreement among these distributions as a signal of epistemic uncertainty, which arises from incomplete knowledge, while consensus suggests more reliable generalization. This can be measured by JSD Cover (1999). Meanwhile, we compute the mean entropy $H$ of individual model predictions to reflect aleatoric uncertainty, capturing inherent input ambiguity. By a novel algorithm that identifies subsets of models $\mathcal{S}$ whose predictions exhibit low disagreement and low intrinsic uncertainty, this formulation enables us to surface high-consensus regions of the input space while balancing the tradeoff between diversity (which may include useful signals) and noise (which can degrade calibration and accuracy).
@@ 4 Multi-LLM selective algorithm
11:  $\hat{p}^{\text{yes}}\leftarrow\text{mean}_{p\in\mathcal{S}}(p^{\text{yes}})$, $u_{\text{total}}\leftarrow u_{\text{epis}}^{\text{prev}}+\beta\cdot u_{\text{alea}}$
@@
Uncertainty Computation. The two types of uncertainty plays an important role in the proposed algorithm. Epistemic uncertainty $\mathcal{U}_{\text{epis}}(\mathcal{S}$ reflects inter-model disagreement and is quantified as the average JSD between each prediction $\mathbf{p}_{i}$ and the subset mean $\bar{\mathbf{p}}$:
@@ 5 Experimental Setup
We use TruthfulQALin et al. (2022), a benchmark of adversarially designed questions with labeled truthful and untruthful answers. The task is to classify each candidate answer as truthful (Yes) or not (No), enabling direct evaluation of both discrimination (AUROC) and calibration (ECE, Brier Score). To evaluate performance in a high-stakes setting, we also apply our method to clinical prediction tasks using two structured EHR datasets: diagnosis prediction from EHRShotWornow et al. (2023) and MIMIC-Extract Wang et al. (2020). On MIMIC, the LLMs predict three outcomes: hospital length of stay $\geq$ 3 days (LOS3), $\geq 7$ days (LOS7), and in-hospital mortality (Mort Hosp.).
We evaluate the following open-source models: Mistral-7B-Instruct Jiang et al. (2023), Gemma-7B-it Team et al. (2023), Qwen2-7B-instruct Yang et al. (2024), and the latest Deepseek-R1-Distil-Qwen-32B (DS-Qwen) DeepSeek-AI (2025). All LLMs are run on a server with 4×A100 40GB GPUs. For DS-Qwen, we apply 8-bit quantization to reduce inference time and memory usage. In addition to single LLM method baseline, we compose two naive multi-LLM baselines: majority voting (counting positive labels), and mean of all LLMs’ $\hat{p}^{yes}$ as the final positive probability.
@@ 6 Results and Discussion
Overall effectiveness. Muse improves both AUROC and calibration metrics compared to single LLMs and naive ensembling baselines, demonstrating its effectiveness in producing reliable and well-calibrated predictions through selective multi-model aggregation. Although the SLL method occasionally yields the highest AUROC, such as DS-Qwen achieving 72.89 on TruthfulQA, it often suffers from poor calibration (ECE 57.30, Brier 54.16). In contrast, Muse offers more balanced predictions, with comparable AUROC (72.35) and substantially lower calibration error (ECE 38.15). Similar gains are observed on the EHRShot acute_mi task, where DS+Mistral+Qwen achieves the best Brier Score (24.4) and strong AUROC (62.2), improving over all single-LLM baselines.
Balancing diversity and noise. A key strength of our approach is its ability to balance diversity with reliability in multi-LLM ensembles. Figure 1 shows that increasing the minimum subset size ($m_{\text{size}}$) and moderately relaxing the epistemic uncertainty threshold ($\epsilon_{\text{tol}}$) consistently improves both AUROC and ECE. A larger $m_{\text{size}}$ ($\geq$ 20) promotes diversity by including more models, while a moderate $\epsilon_{\text{tol}}$ ([0.04, 0.08]) allows controlled disagreement without overwhelming the ensemble with noise. The best performance is achieved when both parameters are carefully balanced. This supports our hypothesis that LLMs offer complementary strengths and provides empirical evidence for our subset-based uncertainty aggregation framework.
Adaptive model selection. Our method further demonstrates adaptive behavior: performance improves when “stronger” LLMs are present but degrades when weak or noisy models dominate the candidate pool, where strength is defined by each model’s single-task performance. This effect is most evident in acute_mi and TQA. In EHRShot, DS and Mistral form a strong ensemble, but adding a weaker model like Gemma introduces noise that contaminates the pool and harms performance. This supports our hypothesis that selective aggregation for trustable consensus is the key to reliable ensemble performance. Muse makes no assumption that more models lead to better results; rather, it selectively aggregates those that contribute meaningful, calibrated signals to reduce total uncertainty. Table 6 includes a pair of contrasting cases from two prediction tasks in EHRShot for readers who’s interested in (“weak models dominate” vs. “encountered strong, stronger”) behavior.
@@ 7 Conclusion
We present Muse, a selective multi-LLM method for uncertainty-aware prediction that improves both accuracy and calibration across domains. Future work will explore dynamic selection strategies and scaling to larger LLM collections in real-world applications, where computational and deployment constraints play a central role.
@@ A.1 Muse Conservative Algorithm
12:  $\hat{p}^{\text{yes}}\leftarrow\text{mean}_{p\in\mathcal{S}}(p^{\text{yes}})$
@@ A.3 More results on MIMIC-Extract
A.4 Analyzing the adapting behavior of Muse method via EHRShot
@@
A.5 Parameter sensitivity on EHRShot

### v2

# Simple Yet Effective: An Information-Theoretic Approach to Multi-LLM Uncertainty Quantification

@@ Abstract
Large language models (LLMs) often behave inconsistently across inputs, indicating uncertainty and motivating the need for its quantification in high-stakes settings. Prior work on calibration and uncertainty quantification often focuses on individual models, overlooking the potential of model diversity. We hypothesize that LLMs make complementary predictions due to differences in training and the Zipfian nature of language, and that aggregating their outputs leads to more reliable uncertainty estimates. To leverage this, we propose MUSE (Multi-LLM Uncertainty via Subset Ensembles), a simple information-theoretic method that uses Jensen-Shannon Divergence to identify and aggregate well-calibrated subsets of LLMs. Experiments on binary prediction tasks demonstrate improved calibration and predictive performance compared to single-model and naïve ensemble baselines. In addition, we explore using MUSE as guided signals with chain-of-thought distillation to fine-tune LLMs for calibration. MUSE is available at:https://github.com/LARK-NLP-Lab/MUSE.
Simple Yet Effective: An Information-Theoretic Approach to Multi-LLM Uncertainty Quantification
Maya Kruse1, Majid Afshar3, Saksham Khatwani1,2, Anoop Mayampurath3, Guanhua Chen3 Yanjun Gao1thanks: Correspondence: yanjun.gao@cuanschutz.edu 1University of Colorado Anschutz Medical Campus 2University of Colorado Boulder 3University of Wisconsin Madison
@@ 1 Introduction
We also ask whether consensus-driven probabilities can teach probabilistic reasoning to individual LLMs. Using Muse outputs as silver-standard supervision for fine-tuning and CoT distillation, we find that ensemble-derived signals are principled but their effectiveness depends on the underlying model, highlighting an important direction for understanding how LLMs internalize probabilistic reasoning.
@@ 3.1 LLM Uncertainty Quantification
To estimate uncertainty, we apply a bootstrapping procedure: we resample $90\%$ of the outputs with replacement and recompute $\hat{p}^{\texttt{yes}}$ over $B=100$ trials (denoting as GENBS). From the resulting $\hat{p}^{\texttt{yes}}$ distribution, we compute variance, entropy, and JSD for our proposed algorithms (§ 3.2).
(2) Sequence Likelihood (SLL) Scoring. We adopt the SLL approach used in prior LLM calibration work. For each input $x$, we compute the total log-likelihood of two candidate completions: "Answer is Yes" and "Answer is No", denoted $\text{LL}_{\texttt{yes}}$ and $\text{LL}_{\texttt{no}}$. These are computed using left-to-right autoregressive decoding: $\text{LL}_{\texttt{label}}=\sum_{t=1}^{T}\log P(y_{t}^{\texttt{label}}\mid x,y_{<t}^{\texttt{label}})$. The final prediction probability is obtained via softmax normalization. We use this predicted distribution to compute AUROC, ECE and Brier Scores Guo et al. (2017). Although not robust to LLM output variability, SLL provides a deterministic scoring baseline for comparison.
3.2 Multi-LLM selective algorithm
The central idea of Muse is that disagreement among LLM predictive distributions signals epistemic uncertainty, while consensus indicates more reliable generalization. This can be measured by JSD Cover (1999). Meanwhile, we compute the mean entropy $H$ of individual model predictions to reflect aleatoric uncertainty, capturing inherent input ambiguity. By a novel algorithm that identifies subsets of models $\mathcal{S}$ whose predictions exhibit low disagreement and low intrinsic uncertainty, this formulation enables us to surface high-consensus regions of the input space while balancing the tradeoff between diversity (which may include useful signals) and noise (which can degrade calibration and accuracy).
@@ 3.2 Multi-LLM selective algorithm
[…]
@@
Uncertainty Computation. The two types of uncertainty plays an important role in the proposed algorithm. Epistemic uncertainty $\mathcal{U}_{\text{epis}}(\mathcal{S})$ reflects inter-model disagreement and is quantified as the average JSD between each prediction $\mathbf{p}_{i}$ and the subset mean $\bar{\mathbf{p}}$:
@@
3.3 Uncertainty-Aware Supervised Fine-Tuning
We test whether Muse-derived probabilities improve model accuracy and calibration through supervised fine-tuning (SFT). We evaluate SFT variants that differ in how probabilities are injected into prompts and whether explicit reasoning is included.
Direct SFT.
We inject probabilistic signals via two prompt formats: (i) Default: include the Muse consensus $\hat{p}$ with the patient input and gold label; (ii) RawProb: replace $\hat{p}$ with bootstrapped per-model probabilities (e.g., Mistral: [0.62, 0.64, 0.60]; Qwen: [0.58, 0.55, 0.57]; DS-Qwen: [0.55, 0.53, 0.56]).
Chain-of-thought distillation.
We optionally append teacher-generated reasoning using three variants: (i) Original (assess whether $\hat{p}$ is reasonable; Table 10), (ii) Bayesian (explicit Bayes framing), and (iii) No $\hat{p}$ (reason without $\hat{p}$). Each CoT variant is paired with both SFT formats.
@@ 4 Experimental Setup
We use TruthfulQALin et al. (2022), a benchmark of adversarially designed questions with labeled truthful and untruthful answers. The task is to classify each candidate answer as truthful (Yes) or not (No), enabling direct evaluation of both discrimination (AUROC) and calibration (ECE, Brier Score). We also apply our method to clinical prediction tasks using two structured EHR datasets: diagnosis prediction from EHRShotWornow et al. (2023) and MIMIC-Extract Wang et al. (2020). On MIMIC, the LLMs predict three outcomes: hospital length of stay $\geq$ 3 days (LOS3), $\geq 7$ days (LOS7), and in-hospital mortality (Mort Hosp.).
We evaluate the following open-source models: Mistral-7B-Instruct Jiang et al. (2023), Gemma-7B-it Team et al. (2023), Qwen2-7B-instruct Yang et al. (2024), and the latest Deepseek-R1-Distil-Qwen-32B (DS-Qwen) DeepSeek-AI (2025). All LLMs are run on a server with 4×A100 40GB GPUs. For DS-Qwen, we apply 8-bit quantization to reduce inference time and memory usage. In addition to single LLM method baseline, we compose two naive multi-LLM baselines: majority voting (counting positive labels), and mean of all LLMs’ $\hat{p}^{yes}$ as the final positive probability. On the Uncertainty-aware SFT experiments, we investigate on two models, Mistral-7B-Instruct and Qwen2-7B-instruct, using the MIMIC-Extract dataset for the length-of-stay 3 days (LOS3) prediction task (y=1y{=}1 if LOS $\geq 3$ days, 0 otherwise). For CoT distillation, we leverage a HIPAA-Compliant Microsoft Azure GPT-o3-mini as the teacher model, and generate 300 samples on LOS-3 training set for each setting. This Azure GPT instances provides high-quality rationales while remaining compliant with the MIMIC-III data use agreement.
MUSE Task
Total
GEM
Qwen
DS-Qwen
Mistral
Conserv LOS 3 31.6 56.97 35.67 25.01 17.42
Greedy LOS 3 90.58 27.84 26.76 25.77 27.01
Conserv TQA 33.33 64.44 19.79 7.89 28.90
Greedy TQA 76.18 33.46 26.89 17.17 31.95
Method Total% GEM% Qwen% DS-Qwen%
Greedy 73.13 47.95 39.90 24.71
Conserv. 53.53 69.16 30.69 13.03
Table 4: Breakdown of LLM inclusion in Muse. Top: frequency of each model’s selection under Conservative and Greedy variants on LOS 3 and TruthfulQA (TQA). Bottom: inclusion frequencies when Mistral is removed from the model pool. (GEM: Gemma-7B, Qwen: Qwen2, DS-Qwen: Deepseek-distilled-Qwen32B, Mistral: Mistral-7B-Instruct).
[…]
@@ 5 Results and Discussion
We organize results around two questions. Q1: Is Muse effective for uncertainty estimation? We evaluate calibration and robustness on general and clinical datasets. Q2: Can Muse provide silver-standard supervision for probabilistic reasoning? We test whether consensus-derived probabilities aid SFT and CoT distillation. Findings show Muse improves calibration and robustness (Q1), while supervision results are mixed: some models benefit, others do not. Overall, multi-LLM aggregation is promising for uncertainty estimation but presents challenges to improve single-model reasoning.
5.1 MUSE Effectiveness
Muse improves both AUROC and calibration metrics compared to single LLMs and naive ensembling baselines, demonstrating its effectiveness in producing reliable and well-calibrated predictions through selective multi-model aggregation. Although the SLL method occasionally yields the highest AUROC, such as DS-Qwen achieving 72.89 on TruthfulQA, it often suffers from poor calibration (ECE 57.30, Brier 54.16). In contrast, Muse offers more balanced predictions, with comparable AUROC (72.35) and substantially lower calibration error (ECE 38.15). Similar gains are observed on the EHRShot acute_mi task, where DS+Mistral+Qwen achieves the best Brier Score (24.4) and strong AUROC (62.2), improving over all single-LLM baselines.
Demonstrating MUSE Impact. We address two-subquestions that illustrate the practical significance of the MUSE algorithm.
(1) Does Muse blindly include weaker models? Table 4 reports selection frequencies under MUSE (Conservative, Greedy). Mistral is chosen less often than Greedy, only when it improves subset consistency. Even strong models (e.g., Qwen) are not always selected, indicating instance-specific, uncertainty-driven choice rather than global accuracy. When a weaker model adds noise, MUSE adapts. Removing Mistral shifts weight to GEM and Qwen, confirming adaptive selection rather than identity-based filtering.
(2) Are gains just from excluding a bad model? Ablations over fixed subsets (Tables 1, 2) and exhaustive pairwise/three-way combinations (Table 8) show that while naive ensembles (e.g., DS+Qwen2) can perform well, MUSE often matches or exceeds them, even when weaker models remain, confirming that it performs input-level rather than fixed global selection. Additional analysis of per-model divergence from the consensus (Table 9) shows small JSD values (<0.1<0.1), supporting our claim that ensemble diversity with selective aggregation underlies MUSE’s effectiveness.
5.2 MUSE-Guided Supervised Fine-Tuning
Figure 1 summarizes the impact of incorporating MUSE-derived probabilities into supervised fine-tuning. Direct SFT yields modest, model-dependent effects. For Mistral, $\hat{p}$ sometimes improves discrimination but often worsens calibration, while GEN consistently lowers ECE. Qwen shows the opposite: calibration improves, but AUROC gains are limited or negative. Adding CoT distillation introduces wider variation. Mistral sees large AUROC increases with RawProb, though often at the cost of calibration. Qwen benefits most from Bayesian CoT, with clear calibration gains in GEN. The results highlight that Muse-derived uncertainty can provide useful supervision, but effects depend on model, supervision style, and whether signals are contextualized through reasoning.
@@ 6 Conclusion
We present Muse, a multi-LLM framework for uncertainty estimation that aggregates predictive distributions into calibrated, uncertainty-aware outputs. Beyond improving accuracy and calibration, we explored using Muse as a supervisory signal for fine-tuning single LLMs, with mixed but promising results, indicating a direction to explore further.
Acknowledgments
This work is supported by U.S. National Library of Medicine R00 LM014308.
@@ A.1 Muse Conservative Algorithm
[…]
@@
Model on LOS3 JSD (Conserv) JSD (Greedy)
GEM 0.0444 0.0709
Qwen2 0.0971 0.0532
DS-Qwen 0.0640 0.06512
Mistral 0.0948 0.05126
Table 9: Comparison of JSD between individual LLM and final MUSE predicted probability. LLM notation is consistent with table LABEL:tab:muse_inclusion
Base CoT Prompt: Generate a short chain-of-thought (CoT) reasoning paragraph (maximum 200 words) that explains the significance of the given p_hat value in the context of hospital length of stay (LOS) prediction. Bayesian CoT Prompt: Generate a short chain-of-thought (CoT) reasoning paragraph (maximum 200 words) that explains the significance of the given p_hat value in the context of hospital length of stay (LOS) prediction. Use Bayesian reasoning to justify if p_hat is a reasonable estimate. If p_hat aligns with the true label, explain why it succeeds. If it does not align, explain why it fails. Additional Information (for both prompts): – The input field contains clinical data from the MIMIC dataset. – The p_hat value is the predicted probability of the patient staying three or more days. – The y_true is the ground-truth label (0 = LOS <3<3 days, 1 = LOS $\geq 3$ days).
Table 10: Base and Bayesian chain-of-thought prompts used for supervised fine-tuning.
With $\hat{p}$ No $\hat{p}$
[…]
Mistral default SLL 54.83 39.51 31.84 47.82 61.17 45.69
GENBS 49.28 29.39 20.00 50.59 38.82 29.75
RP SLL 43.46 53.75 54.04 41.48 59.28 59.28
GENBS 47.76 29.19 16.63 52.08 39.72 31.49
Qwen default SLL 49.52 58.89 46.23 48.56 45.25 44.34
GENBS 44.31 49.97 20.70 51.58 37.43 15.82
RP SLL 55.35 38.94 36.99 51.00 46.54 44.77
GENBS 51.86 41.52 15.29 48.20 37.36 18.19
Table 11: Direct SFT on Mistral-7B and Qwen2-7B under default and RawProb (RP) settings, with and without inclusion of $\hat{p}$. Bold indicates improvement over the no-SFT baselines: Mistral-7B: SLL 41.41/11.34/26.21, GENBS 51.28/15.63/28.37; Qwen2-7B: SLL 56.40/36.50/35.96, GENBS 56.80/31.10/34.18, reported in Tab 3 and Tab 6.
Original Bayesian No $\hat{p}$
Model SFT Output AUROC Brier ECE AUROC Brier ECE AUROC Brier ECE
Mistral default SLL 54.22 42.90 42.97 42.14 43.04 41.03 39.27 35.09 31.23
GENBS 51.91 27.97 14.66 49.30 28.80 16.48 48.52 41.73 31.03
RawProb SLL 44.79 57.75 57.78 58.73 28.90 21.88 64.51 51.46 52.80
GENBS 49.16 28.97 17.60 47.56 29.66 17.38 52.05 42.01 33.29
Qwen default SLL 50.34 52.56 52.08 56.35 30.76 23.23 48.89 36.79 32.53
GENBS 49.72 44.55 16.63 49.77 43.65 16.49 52.95 37.29 17.59
RawProb SLL 48.07 36.16 32.97 45.43 58.95 59.01 43.63 51.21 49.74
GENBS 52.55 40.98 15.19 55.72 39.83 13.81 47.94 40.25 19.23
Table 12: SFT on Mistral-7B and Qwen2-7B with CoT prompts (Original, Bayesian, No $\hat{p}$), under Default/RawProb CoT and SLL/GENBS. Bold values indicate improvements over the no-SFT baselines (Qwen2-7B: SLL 56.40/36.50/35.96, GENBS 56.80/31.10/34.18; Mistral-7B: SLL 41.41/11.34/26.21; GENBS 51.28/15.63/28.37), reported in Tab 3 and Tab 6.
A.2 Balancing diversity and noise.
A key strength of our approach is its ability to balance diversity with reliability in multi-LLM ensembles. Figure 2 shows that increasing the minimum subset size ($m_{\text{size}}$) and moderately relaxing the epistemic uncertainty threshold ($\epsilon_{\text{tol}}$) consistently improves both AUROC and ECE. A larger $m_{\text{size}}$ ($\geq$ 20) promotes diversity by including more models, while a moderate $\epsilon_{\text{tol}}$ ([0.04, 0.08]) allows controlled disagreement without overwhelming the ensemble with noise. The best performance is achieved when both parameters are carefully balanced. This supports our hypothesis that LLMs offer complementary strengths and provides empirical evidence for our subset-based uncertainty aggregation framework.
A.3 Adaptive model selection.
Our method further demonstrates adaptive behavior: performance improves when “stronger” LLMs are present but degrades when weak or noisy models dominate the candidate pool, where strength is defined by each model’s single-task performance. This effect is most evident in acute_mi and TQA. In EHRShot, DS and Mistral form a strong ensemble, but adding a weaker model like Gemma introduces noise that contaminates the pool and harms performance. This supports our hypothesis that selective aggregation for trustable consensus is the key to reliable ensemble performance. Muse makes no assumption that more models lead to better results; rather, it selectively aggregates those that contribute meaningful, calibrated signals to reduce total uncertainty. Table 7 includes a pair of contrasting cases from two prediction tasks in EHRShot for readers who’s interested in (“weak models dominate” vs. “encountered strong, stronger”) behavior.
A.4 Comparison of P(Yes) vs. $U(\mathcal{S})$
A.5 More results on MIMIC-Extract
@@
A.6 Muse JSD from the four LLMs
Table 9 reports the average JSD between each LLM and the Muse consensus. Both weaker and stronger models show similarly small divergences, with no clear relation between a model’s standalone performance and its JSD to MUSE.
A.7 Analyzing the adapting behavior of Muse method via EHRShot
@@
A.8 Parameter sensitivity on EHRShot
@@
A.9 Prompting Strategy for Chain-of-Thought
Table 10 shows the prompts used for generating the base and Bayesian chain-of-thought reasoning paragraphs.
A.10 Detailed results of MUSE-guided SFT
When comparing against the no-SFT baselines, Direct SFT shows mixed improvements. For Mistral, SLL with $\hat{p}$ markedly improves AUROC (41.41 $\to$ 54.83), though calibration deteriorates (ECE 26.21 $\to$ 31.84). GENBS settings yield the strongest calibration benefits, lowering ECE to 16.63 compared to the baseline 28.37, while maintaining competitive AUROC (52.08). For Qwen, SLL with RawProb achieves AUROC (55.35) close to baseline (56.40) but with worse calibration (ECE 36.99 vs. 34.18). By contrast, GENBS achieves notable calibration gains, reducing ECE to 15.29–18.19 compared to the baseline 34.18, though AUROC remains lower. Overall, Direct SFT demonstrates that introducing $\hat{p}$ or raw probability signals can help calibration, especially under GENBS, but often at the cost of discrimination.
CoT distillation SFT (Table 12).
Adding teacher-generated reasoning further diversifies the outcomes. For Mistral, Bayesian RawProb SLL produces the highest AUROC (58.73 vs. baseline 41.41), while No $\hat{p}$ RawProb reaches 64.51, albeit with very poor calibration (ECE 52.80). GENBS again proves more stable, with Original CoT reducing ECE to 14.66, and Bayesian CoT to 16.48, compared to the baseline 28.37. For Qwen, Bayesian Default SLL achieves balanced improvement, raising AUROC to 56.35 while reducing ECE to 23.23 (baseline 34.18). Similarly, GENBS with RawProb Bayesian CoT yields the lowest calibration error overall (ECE 13.81). These findings suggest that CoT distillation can help models internalize probabilistic reasoning and achieve strong calibration improvements, though sometimes at the expense of discrimination (e.g., Qwen SLL RawProb).
The results highlight that Muse-derived uncertainty can provide useful supervision, but the benefits are strongly dependent on LLM, supervision style, and whether the probabilistic signals are contextualized through reasoning.